anybody help me with this 2 questions...having headache....
a ahip traveling east at 25mph is 10mi from a harbor when another ship leaves the harbor traveling east at 35mph. How long does it take the second ship to catch up to the first ship?
Two cyclist start at the same time from opposite ends of a course that is 51mi long. one cyclist is riding at a rate of 16mph and the second cyclist is riding at a rate of 18mph. How long after they begin will they meet?

Answer by mananth(16949)

Ship I speed 25 mph
Ship II speed 35 mph
catch up distance = 10 miles
catchup speed = 35 -25 = 10 mph
catchup time = catchup distance / catch up speed
10/10
=1 hour to catch up
........
cyclist I speed - 16 mph
cyclist II speed =18 mph
riding towards each other.
combined speed = 34 mph
distance = 51 miles
t=d/r
51/34
=1.5 hours